The product of two consecutive integers is 17 less than the square of the larger integer. Find the integers.
Answer by lwsshak3(11628)

The product of two consecutive integers is 17 less than the square of the larger integer. Find the integers.
***
let x=smaller consecutive integer
x+1=larger consecutive integer
..
x(x+1)=(x+1)^2-17
x^2+x=x^2+2x+1-17
x=16
x+1=17
smaller consecutive integer=16
larger consecutive integer=17